# On Geometry-Enhanced Parameter-Efficient Fine-Tuning for 3D Scene Segmentation

### v1

@@ Abstract
The emergence of large-scale pre-trained point cloud models has significantly advanced 3D scene understanding, but adapting these models to specific downstream tasks typically demands full fine-tuning, incurring high computational and storage costs. Parameter-efficient fine-tuning (PEFT) techniques, successful in natural language processing and 2D vision tasks, would underperform when naively applied to 3D point cloud models due to significant geometric and spatial distribution shifts. Existing PEFT methods commonly treat points as orderless tokens, neglecting important local spatial structures and global geometric contexts in 3D modeling. To bridge this gap, we introduce the Geometric Encoding Mixer (GEM), a novel geometry-aware PEFT module specifically designed for 3D point cloud transformers. GEM explicitly integrates fine-grained local positional encodings with a lightweight latent attention mechanism to capture comprehensive global context, thereby effectively addressing the spatial and geometric distribution mismatch. Extensive experiments demonstrate that GEM achieves performance comparable to or sometimes even exceeding full fine-tuning, while only updating 1.6% of the model’s parameters, fewer than other PEFT methods. With significantly reduced training time and memory requirements, our approach thus sets a new benchmark for efficient, scalable, and geometry-aware fine-tuning of large-scale 3D point cloud models. Code will be released.
Figure 1: (a) Existing PEFT methods, such as adapters, prompt tuning, and LoRA, focus on adaptations inside attention and feed-forward layers. (b) In contrast, Geometry Encoding Mixer (GEM) explicitly encodes the geometric cues and mixes them into the pre-trained model, by the Spatial Adapter refining the pre-trained positional encoding and the Context Adapter complementing the local attention. (c) By capturing 3D spatial details and scene-wide geometry context, GEM reaches full fine-tuning performance while tuning $<2\%$ parameters, outperforming existing PEFT methods.
@@ 1 Introduction
Point cloud semantic segmentation is a fundamental task for scene understanding that underpins many real-world applications, from autonomous driving and unmanned aerial vehicles to augmented reality [20, 81, 46]. Recent breakthroughs in large-scale models for language and 2D vision [16, 50, 58, 14, 57, 49] have spurred an analogous trend toward training powerful point cloud backbones capable of capturing rich semantics from 3D scenes [74, 75, 77].
Despite significant capabilities, adapting large pre-trained backbone models to specific downstream target domains is non-trivial. The typical strategy heavily relies on full fine-tuning, a process that is both computationally expensive and storage-intensive, since each downstream task demands an independent copy of all model parameters. Furthermore, fine-tuning large models typically requires extensive datasets to prevent overfitting and catastrophic forgetting [3, 51], while large datasets are usually inaccessible when adapted to smaller, specialized tasks. To address these issues, parameter-efficient fine-tuning (PEFT) methods offer a promising performance in reducing the cost of adaptation while reaching the performance of full fine-tuning with significantly fewer parameters.
While PEFT has been extensively validated in language processing and 2D vision tasks, its effectiveness on large-scale 3D point clouds remains largely unexplored and inadequately addressed. As demonstrated empirically in Fig. 1, existing PEFT methods only yield limited performance gains if directly applying them to a 3D point cloud backbone. We tend to attribute this limitation to a key challenge that is often overlooked by current PEFT methods when using for 3D point cloud: unlike structured data such as text or images, point clouds are inherently unordered sets of coordinates in $\mathbb{R}^{3}$. They exhibit strong irregularity, sparsity, and structural variability, shaped by different sensing protocols and scene geometries. These factors lead to significant geometric and spatial distribution shifts between large-scale pre-training datasets and downstream domains, which existing PEFT methods fail to account for. In particular, representative PEFT methods, such as LoRA [26], adapters [25] and prompt tuning [37, 36], either adapt models at an isolated, per-point level or insert fixed external tokens at the global level, thus failing to adequately capture important geometric and spatial contexts inherent in 3D scenes. Moreover, to avoid the prohibitive computational cost of global attention over a large number of points, current 3D transformers predominantly employ local attention mechanisms without explicitly modeling global contexts, which further restricts the potential of current PEFT approaches in performance.
@@
Empirically, we validate our approach on large-scale 3D scene datasets, including both indoor [12, 60, 1, 82] and outdoor [4] scenes. Our results indicate that models equipped with GEM consistently achieve performance matching or sometimes surpassing full fine-tuning methods, updating merely ~1.6% of model parameters, while being fewer than existing PEFT alternatives. These findings underscore the importance of explicitly modeling geometric and spatial contexts for efficient and effective adaptation in 3D scene understanding. To the best of our knowledge, our work represents the first exploration and validation of PEFT approaches tailored explicitly for large point cloud transformer under large-scale 3D scenes, hoping to establish a foundation for future research and practical deployments.
@@ 2 Related Work
Point clouds serve as an efficient representation for large-scale 3D scenes. Consequently, point cloud segmentation becomes a fundamental task that drives the design of 3D backbone architectures. Since the seminal work PointNet [53, 54], numerous 3D backbone architectures have been proposed for this task. These backbones either project the points onto a grid-like structure, e.g. 3D voxel grid, to exploit 3D convolutional networks [10, 69, 64, 21, 19, 13], or directly process the raw unordered point sets [71, 38, 73, 27, 65, 45]. Recently, inspired by the success of large transformer models in natural language and 2D vision [67, 16, 6, 9], researchers have started training increasingly large and powerful point cloud backbones [55, 88, 68, 56, 34, 76, 75]. Although these models achieve strong performance, they typically need to be trained from scratch for each new dataset and often require specialized training recipes to reach optimal results.
Following the paradigm of large-scale pre-training that has proven successful in vision and language domains [58, 33, 50, 14, 49, 23], researchers have begun exploring similar strategies for 3D point clouds. In particular, self-supervised learning (SSL) on large 3D scene datasets has demonstrated promising results [80, 44, 24, 78]. However, most such efforts focus on modest convolutional backbones like SparseUNet [10] rather than transformer-based architectures. Only recently has an SSL approach been applied to a transformer-based 3D backbone: Sonata [74] introduced SSL for a large point transformer encoder [75], but even this method still requires full fine-tuning on each downstream dataset to achieve optimal performance.
@@
The ever-growing size of transformer-based foundation models[16, 33, 66] makes full fine-tuning for downstream task prohibitively expensive in both memory and computation. PEFT methods address this challenge by adapting large models while updating only a small fraction of their parameters. Existing approaches fall into four broad categories.
Selective fine-tuning methods update a carefully chosen subset of the original weights. One simple variant is linear probing [50, 15] that trains only the classification head. Other strategies restrict training to specific parts of the network, e.g., tuning only bias terms [5] and selecting a subset of parameters based on gradient magnitude criteria [87].
Adapter-based tuning inserts lightweight bottleneck modules into an otherwise frozen backbone. First explored for CNNs [59] and later extended to transformers, adapters may be placed sequentially [25, 61] or in parallel [22, 90, 17, 8] to the original modules. Beyond single task, adapters can be composed or fused, promoting knowledge sharing without catastrophic forgetting [70, 51].
Prompt-based tuning techniques, including prompt tuning [36] and prefix tuning [37], extend to the prompting paradigm in large language models [41]. Instead of hand-crafted prompts, these methods learn task-specific vectors that are prepended to the input sequence [2, 36] or injected as extra tokens in transformer layers [37, 43, 31], thereby steering the model without modifying its original weights [36, 41].
Low-rank adaptation (LoRA) constrains the weight updates to a low-dimensional subspace, by learning a pair of low-rank matrices that are added to each pre-trained weight, typically in the attention layers [26]. This design approximates the effect of full fine-tuning while introducing only a small number of additional parameters. Successors further enhance the approximation ability, robustness, or rank allocation [32, 42, 84, 79, 40].
Recent works also propose hybrid schemes that highlight common design principles [28], for instance, integrating their respective strengths under a unified adapter framework [22].
@@
In the context of 3D point clouds, PEFT remains relatively underexplored. Existing methods thus far have focused mostly on object-level inputs with limited spatial scale. Some approaches introduce prompt tokens that adapt to 3D data, where the prefix tokens are dynamically generated from intermediate features [83, 89] or from the spatial centers of local patches [85]. Several works also introduce auxiliary graph-based side networks, operating either in the spectral domain [39] or the spatial domain [18], to provide global geometric context. Other methods construct banks of prompt vectors using domain-specific dataset to inject 3D prior knowledge [63], or combine language-side prompt tuning with point-cloud adapters to handle open-vocabulary recognition [62].
Compared to object-level datasets [7, 48], 3D scene understanding involves much larger inputs containing on the order of millions of points [12, 1]. This scale amplifies computational challenges and underscores the need for PEFT techniques expressly designed for large 3D scene models, which is however a research direction that remains largely unexplored.
@@ 3 Methodology
We propose the Geometry Encoding Mixer (GEM) as a lightweight parameter-efficient module for fine-tuning point-cloud transformer. Fig. 2 summarises the overall architecture.
In particular, GEM consists of two complementary adaptations: a Spatial Adapter that refine the positional encoding of each point, injecting local geometry information overlooked by generic adapters; and a Context Adapter that distills a compact set of latent tokens that broadcast global scene-level cues. Both components follow the residual, bottleneck design of classic adapters, yet emphasize and operate on spatial rather than channel space.
@@ 3.1 Preliminaries
Chanlleges under 3D.
Following transformer [67, 16], point-cloud transformers also build upon self-attention layer (Attn) followed by feed-forward network (FFN). However, due to the large scale input points at million scale, global self-attention over all points would demand prohibitively large GPU memory and compute resource, if ever possible, due to the quadratic complexity of the attention operations: $\mathcal{O}(\text{Attn})=n^{2}$, where $n$ denotes the number of input points. In this regards, state-of-the-art point-cloud transformers propose various design of local attentions [88, 76, 75, 34, 68] to cap the complexity, where each point can only attend to points within the same patch. With a patch size of $p$, the complexity of local attention is then $\mathcal{O}(\text{Attn}_{\text{loc}})=np$. For example, it is set to $p=16$ in PT [88] and $p=1024$ in the larger PTv3 [75].
The irregular and sparse nature of 3D point cloud, together with these architecture design, greatly imped the application of standard PEFT methods such as adapter, LoRA and prompt tuning. As shown in Fig. 1, directly applying existing methods yields only marginal gains. We attribute this to their lack of spatial structure awareness of 3D geometry and the inability to capture scene-wide context under the local attention regimes of modern point cloud backbones.
@@
The adapter-based methods [25, 61] insert small modules after the attention or FFN layers:
@@
where $\bm{W}_{\text{down}}\in\mathbb{R}^{d\times r}$ and $\bm{W}_{\text{up}}\in\mathbb{R}^{r\times d}$ are down and up projections that forms a bottleneck structure with $r\ll d$ and activation $\sigma$ in the mid, and are also surrounded by residual connection. As generic MLP bottlenecks, adapters adapt the model on a per-point basis, overlooking the important spatial cues in 3D point cloud.
@@
LoRA methods [26] typically adapt the attention projection layers and approximate the full fine-tuning with a bottleneck in weights space, transforming the attention into:
$\text{Attn}_{\text{LoRA}}=\text{softmax}\frac{(\bm{Q}+\Delta\bm{Q})(\bm{K}+\Delta\bm{K})^{T}}{\sqrt{d_{k}}}\bm{V},$ (2)
@@
Prompt tuning methods [37, 36] explicitly introduce global tokens into the attention:
$\text{Attn}_{\text{prompt}}=\text{softmax}\frac{\bm{Q}[\bm{P}_{K};\bm{K}]^{T}}{\sqrt{d_{k}}}[\bm{P}_{V};\bm{V}],$ (3)
where $\bm{P}_{K},\bm{P}_{V}\in\mathbb{R}^{m\times d}$ are two set of tokens prepending to the original key-value pair. In spite of the constraint of local patches, prompt tokens can be duplicated into each patch. However, with $m<<n$, it can be hard for a few static external tokens to capture scene-specific context that can vary across point clouds within the same dataset. In addition, it also overlooks the spatial pattern of the point cloud, lacking the adaptation on local point features.
@@
There are more methods that explore other aspects. Selective tunings propose to tune a small subset of the forzen model parameters, such as BitFit [5] that tunes only bias terms. Hybrid methods are also explored, such as MAM [22] that couples FFN adapter with prompt tuning for attention layer. These methods yet still inherit the constraints from those representative PEFT methods, which are hindered by the local attention and fails to recognize the spatial structure.

### v2

@@ Abstract
The emergence of large-scale pre-trained point cloud models has significantly advanced 3D scene understanding, but adapting these models to specific downstream tasks typically demands full fine-tuning, incurring high computational and storage costs. Parameter-efficient fine-tuning (PEFT) techniques, successful in natural language processing and 2D vision tasks, would underperform when naively applied to 3D point cloud models due to significant geometric and spatial distribution shifts. Existing PEFT methods commonly treat points as orderless tokens, neglecting important local spatial structures and global geometric contexts in 3D modeling. To bridge this gap, we introduce the Geometric Encoding Mixer (GEM), a novel geometry-aware PEFT module specifically designed for 3D point cloud transformers. GEM explicitly integrates fine-grained local positional encodings with a lightweight latent attention mechanism to capture comprehensive global context, thereby effectively addressing the spatial and geometric distribution mismatch. Extensive experiments demonstrate that GEM achieves performance comparable to or sometimes even exceeding full fine-tuning, while only updating 1.6% of the model’s parameters, fewer than other PEFT methods. With significantly reduced training time and memory requirements, our approach thus sets a new benchmark for efficient, scalable, and geometry-aware fine-tuning of large-scale 3D point cloud models. Code is available at https://github.com/LiyaoTang/GEM.
11footnotetext: Corresponding authors.11footnotetext: Also affiliated with Cisco - La Trobe Centre for AI and IoT.22footnotetext: College of Computing and Data Science, NTU, Singapore 639798.
Figure 1: (a) Existing PEFT methods, such as adapters, prompt tuning, and LoRA, focus on adaptations inside attention and feed-forward layers. (b) In contrast, Geometry Encoding Mixer (GEM) explicitly encodes the geometric cues and mixes them into the pre-trained model, by the Spatial Adapter refining the pre-trained positional encoding and the Context Adapter complementing the local attention. (c) By capturing 3D spatial details and scene-wide geometry context, GEM reaches full fine-tuning performance while tuning $2\%$ parameters, outperforming existing PEFT methods.
@@ 1 Introduction
Point cloud semantic segmentation is a fundamental task for scene understanding that underpins many real-world applications, from autonomous driving and unmanned aerial vehicles to augmented reality [ptsurvey, ptsurvey_seg, ptsurvey_tr]. Recent breakthroughs in large-scale models for language and 2D vision [vit, dinov2, clip, deepseekv3, qwen25, gpt4o] have spurred an analogous trend toward training powerful point cloud backbones capable of capturing rich semantics from 3D scenes [sonata, ptv3, ppt].
Despite significant capabilities, adapting large pre-trained backbone models to specific downstream target domains is non-trivial. The typical strategy heavily relies on full fine-tuning, a process that is both computationally expensive and storage-intensive, since each downstream task demands an independent copy of all model parameters. Furthermore, fine-tuning large models typically requires extensive datasets to prevent overfitting and catastrophic forgetting [adap_simple, adap_fusion], while large datasets are usually inaccessible when adapted to smaller, specialized tasks. To address these issues, parameter-efficient fine-tuning (PEFT) methods offer a promising performance in reducing the cost of adaptation while reaching the performance of full fine-tuning with significantly fewer parameters.
While PEFT has been extensively validated in language processing and 2D vision tasks, its effectiveness on large-scale 3D point clouds remains largely unexplored and inadequately addressed. As demonstrated empirically in Fig.˜1, existing PEFT methods only yield limited performance gains if directly applied to a 3D point cloud backbone. We tend to attribute this limitation to a key challenge that is often overlooked by current PEFT methods when using for 3D point cloud: unlike structured data such as text or images, point clouds are inherently unordered sets of coordinates in $\mathbb{R}^{3}$. They exhibit strong irregularity, sparsity, and structural variability, shaped by different sensing protocols and scene geometries. These factors lead to significant geometric and spatial distribution shifts between large-scale pre-training datasets and downstream domains, which existing PEFT methods fail to account for. In particular, representative PEFT methods, such as LoRA [lora], adapters [adap] and prompt tuning [prefix, prefix_prompt], either adapt models at an isolated, per-point level or insert fixed external tokens at the global level, thus failing to adequately capture important geometric and spatial contexts inherent in 3D scenes. Moreover, to avoid the prohibitive computational cost of global attention over a large number of points, current 3D transformers predominantly employ local attention mechanisms without explicitly modeling global contexts, which further restricts the potential of current PEFT approaches in performance.
@@
Empirically, we validate our approach on large-scale 3D scene datasets, including both indoor [scannet, scannet200, s3dis, scannetpp] and outdoor [semkitti] scenes. Our results indicate that models equipped with GEM consistently achieve performance matching or sometimes surpassing full fine-tuning methods, updating merely ~1.6% of model parameters, while being fewer than existing PEFT alternatives. These findings underscore the importance of explicitly modeling geometric and spatial contexts for efficient and effective adaptation in 3D scene understanding. To the best of our knowledge, our work represents the first exploration and validation of PEFT approaches tailored explicitly for large point cloud transformer under large-scale 3D scenes, hoping to establish a foundation for future research and practical deployments.
@@ 2 Related Work
Point clouds serve as an efficient representation for large-scale 3D scenes. Consequently, point cloud segmentation becomes a fundamental task that drives the design of 3D backbone architectures. Since the seminal work PointNet [pointnet, pointnet++], numerous 3D backbone architectures have been proposed for this task. These backbones either project the points onto a grid-like structure, e.g. 3D voxel grid, to exploit 3D convolutional networks [minkowski, ocnn, seg_vx_SEGCloud, occuseg, submanifold, seg_vx_ScanComplete], or directly process the raw unordered point sets [dgcnn, pointcnn, pointconv, randlanet, kpconv, closerlook]. Recently, inspired by the success of large transformer models in natural language and 2D vision [transformer, vit, tr_detr, tr_mask2former], researchers have started training increasingly large and powerful point cloud backbones [pointnext, pttransformer, octformer, cdformer, stratified, ptv2, ptv3]. Although these models achieve strong performance, they typically need to be trained from scratch for each new dataset and often require specialized training recipes to reach optimal results.
Following the paradigm of large-scale pre-training that has proven successful in vision and language domains [clip, sam, dinov2, deepseekv3, gpt4o, tr_mae], researchers have begun exploring similar strategies for 3D point clouds. In particular, self-supervised learning (SSL) on large 3D scene datasets has demonstrated promising results [pointcontrast, p4contrast, contrastscene, msc]. However, most such efforts focus on modest convolutional backbones like SparseUNet [minkowski] rather than transformer-based architectures. Only recently has an SSL approach been applied to a transformer-based 3D backbone: Sonata [sonata] introduced SSL for a large point transformer encoder [ptv3], but even this method still requires full fine-tuning on each downstream dataset to achieve optimal performance.
@@
The ever-growing size of transformer-based foundation models[vit, sam, llama] makes full fine-tuning for downstream task prohibitively expensive in both memory and computation. PEFT methods address this challenge by adapting large models while updating only a small fraction of their parameters. Existing approaches fall into four broad categories.
Selective fine-tuning methods update a carefully chosen subset of the original weights. One simple variant is linear probing [dinov2, lin_decaf] that trains only the classification head. Other strategies restrict training to specific parts of the network, e.g., tuning only bias terms [select_bitfit] and selecting a subset of parameters based on gradient magnitude criteria [select_gps].
Adapter-based tuning inserts lightweight bottleneck modules into an otherwise frozen backbone. First explored for CNNs [adap_cnn] and later extended to transformers, adapters may be placed sequentially [adap, adap+] or in parallel [adap_mam, adap_ciat, adap_bias, adap_former] to the original modules. Beyond single task, adapters can be composed or fused, promoting knowledge sharing without catastrophic forgetting [adap_adamix, adap_fusion].
Prompt-based tuning techniques, including prompt tuning [prefix_prompt] and prefix tuning [prefix], extend to the prompting paradigm in large language models [prefix_promptsurvey]. Instead of hand-crafted prompts, these methods learn task-specific vectors that are prepended to the input sequence [prefix_visualprompt, prefix_prompt] or injected as extra tokens in transformer layers [prefix, prefix_promptv2, prefix_vpt], thereby steering the model without modifying its original weights [prefix_prompt, prefix_promptsurvey].
Low-rank adaptation (LoRA) constrains the weight updates to a low-dimensional subspace by learning a pair of low-rank matrices that are added to each pre-trained weight, typically in the attention layers [lora]. This design approximates the effect of full fine-tuning while introducing only a small number of additional parameters. Successors further enhance the approximation ability, robustness, or rank allocation [lora_rs, lora_dora, lora_ada, lora_reft, ia3].
Recent works also propose hybrid schemes that highlight common design principles [adap_llmsurvey], for instance, integrating their respective strengths under a unified adapter framework [adap_mam].
@@
In the context of 3D point clouds, PEFT remains relatively underexplored. Existing methods thus far have focused mostly on object-level inputs with limited spatial scale. Some approaches introduce prompt tokens that adapt to 3D data, where the prefix tokens are dynamically generated from intermediate features [ptpeft_instance, ptpeft_dyadap, ptpeft_gaprompt] or from the spatial centers of local patches [ptpeft_posprompt]. Several works also introduce auxiliary side networks, operating either in the spectral domain [ptpeft_gst] or the spatial domain [ptpeft_stag, ptpeft_pointlora], to provide geometric context. Other methods construct banks of prompt vectors using domain-specific dataset to inject 3D prior knowledge [ptpeft_peft], or combine language-side prompt tuning with point-cloud adapters to handle open-vocabulary recognition [ptpeft_prompt].
Compared to object-level datasets [shapenetpart, partnet], 3D scene understanding involves much larger inputs containing on the order of millions of points [scannet, s3dis]. This scale amplifies computational challenges and underscores the need for PEFT techniques expressly designed for large 3D scene models, which is however a research direction that remains largely unexplored.
@@ 3 Methodology
We propose the Geometry Encoding Mixer (GEM) as a lightweight parameter-efficient module for fine-tuning point-cloud transformer. Fig.˜2 summarises the overall architecture.
In particular, GEM consists of two complementary adaptations: a Spatial Adapter that refines the positional encoding of each point, injecting local geometry information overlooked by generic adapters; and a Context Adapter that distills a compact set of latent tokens that broadcast global scene-level cues. Both components follow the residual, bottleneck design of classic adapters, yet emphasize and operate on spatial rather than channel space.
@@ 3.1 Preliminaries
Challenges under 3D.
Following transformer [transformer, vit], point-cloud transformers also build upon self-attention layer (Attn) followed by feed-forward network (FFN). However, due to the large scale input points at million scale, global self-attention over all points would demand prohibitively large GPU memory and compute resource, if ever possible, due to the quadratic complexity of the attention operations: $\mathcal{O}(\text{Attn})=n^{2}$, where $n$ denotes the number of input points. In this regard, state-of-the-art point-cloud transformers propose various designs of local attentions [pttransformer, ptv2, ptv3, stratified, octformer] to cap the complexity, where each point can only attend to points within the same patch. With a patch size of $p$, the complexity of local attention is then $\mathcal{O}(\text{Attn}_{\text{loc}})=np$. For example, it is set to $p=16$ in PT [pttransformer] and $p=1024$ in the larger PTv3 [ptv3].
The irregular and sparse nature of 3D point cloud, together with these architecture designs, greatly impede the application of standard PEFT methods such as adapter, LoRA, and prompt tuning. As shown in Fig.˜1, directly applying existing methods yields only marginal gains. We attribute this to their lack of spatial structure awareness of 3D geometry and the inability to capture scene-wide context under the local attention regimes of modern point cloud backbones.
@@
The adapter-based methods [adap, adap+] insert small modules after the attention or FFN layers:
@@
where $\bm{W}_{\text{down}}\in\mathbb{R}^{d\times r}$ and $\bm{W}_{\text{up}}\in\mathbb{R}^{r\times d}$ are down and up projections that form a bottleneck structure with $r\ll d$ and activation $\sigma$ in the mid, and are also surrounded by residual connection. As generic MLP bottlenecks, adapters adapt the model on a per-point basis, overlooking the important spatial cues in 3D point cloud.
@@
LoRA methods [lora] typically adapt the attention projection layers and approximate the full fine-tuning with a bottleneck in weight space, transforming the attention into:
$\text{Attn}_{\text{LoRA}}=\text{Attn}(\bm{Q}+\Delta\bm{Q},\bm{K}+\Delta\bm{K},\bm{V}),$ (2)
@@
Prompt tuning methods [prefix, prefix_prompt] explicitly introduce global tokens into the attention:
$\text{Attn}_{\text{prompt}}=\text{Attn}(\bm{Q},[\bm{P}_{K};\bm{K}],[\bm{P}_{V};\bm{V}]),$ (3)
where $\bm{P}_{K},\bm{P}_{V}\in\mathbb{R}^{m\times d}$ are two sets of tokens prepending to the original key-value pair. In spite of the constraint of local patches, prompt tokens can be duplicated into each patch. However, with $m<<n$, it can be hard for a few static external tokens to capture scene-specific context that can vary across point clouds within the same dataset. In addition, it also overlooks the spatial pattern of the point cloud, lacking the adaptation to local point features.
@@
There are more methods that explore other aspects. Selective tunings propose to tune a small subset of the frozen model parameters, such as BitFit [select_bitfit] that tunes only bias terms. Hybrid methods are also explored, such as MAM [adap_mam] that couples FFN adapter with prompt tuning for attention layer. These methods still inherit the constraints from those representative PEFT methods, which are hindered by the local attention and fail to recognize the spatial structure.